# Error Slice Discovery via Manifold Compactness

### v1

@@ 5 Experiments
In this section, we conduct extensive experiments to demonstrate the validity of our proposed metric and the advantages of our algorithm MCSD compared with previous methods. For quantitative results, we conduct experiments on the error slice discovery benchmark dcbench (Eyuboglu et al., 2022). Besides, we conduct experiments on other types of datasets like classification for medical images (Irvin et al., 2019), object detection for driving (Yu et al., 2020), and detection of toxic comments (Borkan et al., 2019), which showcase the great potential of our algorithm to be applied to various tasks. The baselines we compare with are Spotlight (d’Eon et al., 2022), Domino (Hendrycks & Gimpel, 2016), and PlaneSpot (Plumb et al., 2023). More details of baselines are included in Section A.4.
Note that in all our experiments, we apply algorithms to the validation dataset $\{x_{i}^{\text{va}},y_{i}^{\text{va}}\}_{i=1}^{n_{\text{va}}}$ to obtain the slicing function $g_{\varphi}$, and then employ $g_{\varphi}$ on the test dataset $\{x_{i}^{\text{te}},y_{i}^{\text{te}}\}_{i=1}^{n_{\text{te}}}$ to acquire the prediction probability of each test sample belonging to the error slice. We choose the top $\alpha n_{\text{te}}$ samples from the test dataset sorted by the prediction probabilities as the error slice $\hat{\mathcal{S}}$, and calculate evaluation metrics based on it. As for our method that outputs the error slice of the validation dataset instead of a slicing function, to compare with other methods, we additionally train a binary MLP classifier $g_{\varphi}:\mathcal{Z}\mapsto[0,1]$ on top of embeddings, by treating samples in the error slice as positive examples, and treat the rest as negative ones. However, it is worth noting that our method is effective at error slice discovery without this additional slicing function, which is illustrated in Section A.5.
Besides, following previous works of error slice discovery (Eyuboglu et al., 2022; Wang et al., 2023b), for image data, we employ the image encoder of CLIP with a backbone of ViT-B/32 to extract embeddings of images for error slice discovery algorithms in our main experiments. For text data, we employ pretrained BERTbase to extract embeddings. We conduct additional experiments to show that our method is flexible in the choice of the feature extractor $h_{fe}$, whose detailed results are left in Section A.6. Due to the space limit, for all case studies of visual tasks, we only exhibit 3 or 5 images randomly sampled from each identified slice of the test dataset, and for baselines we only exhibit images from the slice identified by Domino, the previous SOTA algorithm. We put more examples including those of Spotlight and PlaneSpot in Section A.7, about 20 images for each identified slice. For running time comparison and related analyses of our method and the baselines, we leave results in Section A.8. For the choice and analyses of hyperparameters, we leave them in Section A.9. For the improvement of the original models utilizing the discovered error slices, we leave results in Section A.10.
In addition to coherence, we also compute the average performance of the given model $f_{\theta_{0}}$ on the identified slice $\hat{\mathcal{S}}$. For classification tasks, the performance metric is average accuracy. For object detection, it could be Average Precision (AP). Note that now there are two evaluation metrics at the same time. In this case, we put more emphasis on coherence instead of performance, since we only require the performance of the identified slice to be low to a certain degree but expect it to be as coherent as possible for the benefits of interpretation. This is similar to dcbench (Eyuboglu et al., 2022) where coherence also outweighs performance and is chosen as the main evaluation metric.
@@
Metric Precision@10 (%) $\uparrow$ Precision@25 (%) $\uparrow$ Average Precision (%) $\uparrow$ Manifold Comp.$\uparrow$
@@ 5.1 Benchmark results: Dcbench
Method Acc. (%) $\downarrow$ Comp.$\uparrow$ Acc. (%) $\downarrow$ Comp.$\uparrow$
@@
Dcbench (Eyuboglu et al., 2022) offers 886 publicly available settings for error slice discovery. Each setting consists of a trained ResNet-18 (He et al., 2016), a validation dataset and a test dataset, both with predefined underperforming slice labels. The validation dataset and its error slice labels are taken as input of slice discovery methods, while the test dataset and its error slice labels are used for evaluation. There are three types of slices in dcbench: correlation slices, rare slices, and noisy label slices. The correlation slices are generated from CelebA (Liu et al., 2015), while the other two types of slices are generated from ImageNet (Deng et al., 2009). More details are included in Section A.11. In terms of evaluation metrics, we employ precision@$k$ and average precision following dcbench’s practice, where precision@$k$ is the proportion of samples with top $k$ highest probabilities output by the learned slicing function that belongs to the predefined underperforming slice, and average precision is calculated based on precision@$k$ with different values of $k$. We also calculate manifold compactness as Equation 1. For all these metrics, a higher value indicates higher coherence of the identified slice, thus implying a more effective algorithm capable of error slice discovery.
Effectiveness of our method
Validity of our metric
It is also worth noting that the proposed metric manifold compactness shows a strong consistency with other metrics. In Table 1, we find that the rank order of the four methods based on precision metrics is always MCSD, Domino, Spotlight, PlaneSpot, which is generally the same as the rank order based on manifold compactness, except for the correlation slice where the rank order of Domino and Spotlight switches. While other metrics require access to labels of predefined underperforming slices, our metric does not rely on them. This demonstrates the validity and advantages of our proposed manifold compactness when measuring coherence and evaluating the error slice discovery algorithms.
5.2 Case Study: CelebA
CelebA (Liu et al., 2015) is a large facial dataset of 202,599 images, each with annotations of 40 binary attributes. In the setting of subpopulation shift, it is the most widely adopted dataset since it is easy to generate spurious correlations between two specific attributes by downsampling the dataset (Yang et al., 2023; Sagawa et al., 2019; Liu et al., 2021a). Different from settings in dcbench, in this case study we follow Sagawa et al. (2019) to treat the binary label of blond hair as the target of prediction and directly use the whole dataset of CelebA (Liu et al., 2015) without downsampling, thus closer to the real scenario. In terms of implementation details, we employ the default data split provided by CelebA and follow the training process of ERM in (Sagawa et al., 2019) to train a ResNet-50. We apply error slice discovery algorithms on both categories respectively, thus taking advantage of outcome labels that are known during slice discovery. We also illustrate results of directly selecting top $\alpha n_{\text{te}}$ samples sorted by prediction losses.
@@ 5.2 Case Study: CelebA
Method Acc. (%) $\downarrow$ Comp.$\uparrow$ Acc. (%) $\downarrow$ Comp.$\uparrow$
@@
From Table 2, we can see that for both categories of CelebA, our algorithm identifies the most coherent underperforming slice in terms of manifold compactness, where higher is better. Although it ranks 2nd for the category of blond hair in terms of accuracy, where lower is better, for the task of error slice discovery, we put more emphasis on coherence since we want the identified slices to be interpretable, and we only require the performance of the slice to be lower than a threshold compared with the overall performance, as stated in Section 3. In Figure 4, left five columns and right five columns are from the two categories separately. The four rows correspond to randomly sampled images from different sources: the error slice that Domino identifies, the error slice that MCSD identifies, top $\alpha n_{\text{te}}$ samples sorted by the loss, and all samples of the corresponding category. We can see that images from the error slice identified by MCSD obviously exhibit more coherent characteristics than others.
@@ 5.3 Case Study: CheXpert
To demonstrate the effectiveness of our algorithm on other types of data, we conduct experiments on a medical imaging dataset, i.e. CheXpert (Irvin et al., 2019), where the task is to predict whether patients are ill or not based on their chest X-ray images. It contains 224,316 images from 65,240 patients. We follow the data split and training process of Yang et al. (2023) to train a ResNet-50. Still, we apply algorithms to images of ill and healthy patients respectively.
@@ 5.4 Case Study: BDD100K
Method AP (%) $\downarrow$ Comp.$\uparrow$ AP (%) $\downarrow$ Comp.$\uparrow$
@@
Compared with most previous algorithms (Eyuboglu et al., 2022; Wang et al., 2023b; Plumb et al., 2023) requiring prediction probabilities as a part of input and are only designed for classification tasks, our algorithm is flexible to be employed in various tasks since it takes prediction losses as input. To illustrate its benefits of extending to other tasks, we conduct a case study on BDD100K (Yu et al., 2020), a large-scale dataset composed of driving scenes with abundant annotations. It includes ten tasks, of which we investigate object detection in our paper. The number of images in its object detection task is 79,863, which we split into train, validation, and test datasets with the ratio 2:1:1. We train a YOLOv7 (Wang et al., 2023a) and try to identify coherent error slices for it. We employ Average Precision (AP) as the performance metric that is widely adopted in detection tasks. Of the 13 categories in the task, we select 2 categories with a relatively high overall performance and a large sample size, i.e. pedestrian and traffic light. We apply our algorithm MCSD for each of them respectively. Note that we could not compare with Domino or PlaneSpot since neither of them is applicable to tasks other than classification.
Table 4 shows that MCSD successfully identifies error slices whose AP are lower than those of overall for both categories, and whose coherence is higher than that of Spotlight in terms of manifold compactness. In Figure 6, each row corresponds to five images randomly sampled from a given source. Left three columns correspond to the category of pedestrians, while right three columns correspond to the category of traffic lights. For both pedestrians and traffic lights, samples from the source of MCSD are coherent in that they are all taken at night. This conforms to the intuition that it is more difficult to recognize and locate objects when the light is poor. However, directly sampling from the high-loss images can hardly exhibit any common patterns. This reveals the potential of MCSD to be extended to other types of tasks.
@@ 5.5 Case Study: CivilComments
In addition to experiments on visual tasks, to demonstrate the applicability of our method to other types of data, we conduct experiments on CivilComments (Borkan et al., 2019), a text dataset of 244,436 comments included in popular distribution shift benchmarks (Yang et al., 2023; Koh et al., 2021). Its task is to predict whether a given comment is toxic or not. We follow the data split and training process of Yang et al. (2023) to train a BERTbase. We apply algorithms to toxic and non-toxic comments respectively. In Table 5, we can see that MCSD identifies slices of the lowest accuracy and highest manifold compactness in both categories.
Table 5: Results on CivilComments, along with the overall accuracy of the trained model. “Comp.” means “Manifold Compactness”. $\uparrow$” indicates that higher is better, while $\downarrow$” indicates that lower is better. We mark the best method in bold type. “%” indicates that the digits are percentage values.
@@
Method Acc. (%) $\downarrow$ Comp.$\uparrow$ Acc. (%) $\downarrow$ Comp.$\uparrow$
@@ 6 Conclusion
In this paper, inspired by the data geometry property, we propose manifold compactness as a metric of coherence given a slice, which does not rely on predefined underperforming slice labels. We conduct empirical analyses to justify the rationality of our proposed metric. With the help of explicit metrics for risk and coherence, we develop an algorithm that directly incorporates both risk and coherence into the optimization objective. We conduct experiments on a benchmark and perform case studies on various types of datasets to demonstrate the validity of our proposed metric and the superiority of our algorithm, along with the potential to be flexibly extended to different types of tasks.
@@ Appendix A Appendix
A.1 More experimental results related to manifold compactness
@@ A.1.1 Visualization Analyses
A.1.2 Comparison with Variance
We compare manifold compactness with variance, a common and representative metric of coherence directly calculated in the Euclidean space, on different semantically predefined slices. Note that since the slices are not of the same size, to compare manifold compactness of different slices properly, for each given slice we randomly sample a subset of size 150 with 20 times, and average the manifold compactness of 20 subsets as the manifold compactness of the given slice. For CelebA, we use the binary label $y$ to indicate whether the person has blond hair or not, and $a$ to indicate whether the person is male or not. From Table 6, we can see that for manifold compactness, its value of the more fine-grained slice, i.e. the more coherent slice, is larger than the more coarse-grained slice. For example, the manifold compactness of $y=1\&a=0$ is $0.38$, larger than that of $y=1$ (the value is $0.35$) or $a=0$ (the value is $0.13$). Such a relationship holds for every pair of slices. However, in terms of variance, for example, variance of $y=1\&a=1$ is $39.6$, larger than that of $y=1$ whose value is $36.2$, which is contrary to our expectation that variance of the more fine-grained slice is smaller than that of the more coarse-grained slice. For CheXpert, we use the binary label $y$ to indicate whether the person is ill or not, and $a$ to indicate whether the person is male or not. We also find that for manifold compactness, its value of the more fine-grained slice, i.e. the more coherent slice, is larger than the more coarse-grained slice, while the value of variance is not consistent with the granularity of the slice. We also additionally compare with other metrics that are directly calculated in Euclidean distance, including Mean Absolute Deviation (MeanAD), Median Absolute Deviation (MedianAD), and Interquartile Range (IQR). We find that they exhibit similar phenomenons to variance, i.e. the metric value of the more coarse-grained slice is sometime even smaller than that of the more fine-grained slice, which contradicts our expectation. Thus we demonstrate that manifold compactness is better at capturing semantic coherence than variance does.
@@ A.3 Theoretical analyses
We initialize $\mathbf{w}^{(0)}=\mathbf{w}^{\prime}$. Then we sweep a variable $i$ from $1$ to $n-1$. For each iteration with $11\leq j\leq n-$ 1, we generate a new weight vector $\mathbf{w}^{(j)}$ by the following process. We can prove that $\mathbf{w}^{(j)}$ is one solution of Equation 2 and $w_{r_{i}}^{(j)}\in\{0,1\},\forall 1\leq i\leq j$ by mathematical induction, which is already satisfied for $j=0$.
Firstly, we assign $w^{(j)}_{r_{i}}=w^{(j-1)}_{r_{i}}$ for $11\leq i\leq j-$ 1 and $j+2\leq i\leq n$ and denote $C=w^{(j-1)}_{r_{j}}+w^{(j-1)}_{r_{j+1}}\in[0,2]$.
@@
A.4 Baselines
The baselines are:
  • Spotlight (d’Eon et al., 2022): It learns a point in the embedding space as the risky centroid, and chooses the closest points to the centroid as the error slice.
  • Domino (Hendrycks & Gimpel, 2016): It develops an error-aware Gaussian mixture model (GMM) by incorporating predictions into the modeling process of GMM.
  • PlaneSpot (Plumb et al., 2023): It combines the prediction confidence and the reduced two-dimensional representation together as the input of a GMM.
A.5 Ablation of the slicing function
In Algorithm 1, after acquiring the desired samples, we additionally train an MLP as the slicing function. This is because the standard evaluation process of error slice discovery, an error slice discovery method is applied to validation data to obtain the slicing function, and then the slicing function is applied to test data to calculate evaluation metrics and conduct case analyses. Such practice is also adopted by dcbench (Eyuboglu et al., 2022), the benchmark of error slice discovery. Thus we follow this practice by additionally train a slicing function after we obtain the desired samples, so that we can compare fairly with previous methods. However, even without training this slicing function, our method can still produce meaningful results of case studies. We show examples randomly sampled from optimized results of Equation 2 in Figure 11 and 12. We can see that there is still high coherence in the identified slices.

### v2

@@ 5 Experiments
In this section, we conduct extensive experiments to demonstrate the validity of our proposed metric and the advantages of our algorithm MCSD compared with previous methods. For quantitative results, we conduct experiments on the error slice discovery benchmark dcbench (Eyuboglu et al. 2022). Besides, we conduct experiments on other types of datasets like classification for medical images (Irvin et al. 2019), object detection for driving (Yu et al. 2020), and detection of toxic comments (Borkan et al. 2019), which showcase the great potential of our algorithm to be applied to various tasks. The baselines we compare with are Spotlight (d’Eon et al. 2022), Domino (Eyuboglu et al. 2022), and PlaneSpot (Plumb et al. 2023). More experimental details are included in Section A.4.
@@
[…]
@@
For evaluation, we compute manifold compactness as the main coherence metric along with the average performance of the given model $f_{\theta_{0}}$ on the identified slice $\hat{\mathcal{S}}$. For classification, the performance metric is accuracy. For object detection, it is Average Precision (AP). Note that there are two aspects of evaluation simultaneously. In this case, we put more emphasis on coherence than performance, since we only require the performance of the identified slice to be low to a certain degree but expect it to be as coherent as possible for the benefits of interpretation. This is similar to dcbench (Eyuboglu et al. 2022) where coherence outweighs performance and is chosen as the main evaluation metric.
For running time comparison and related analyses of our method and the baselines, we leave results in Section A.5. For the choice and analyses of hyperparameters, we leave them in Section A.6. For the improvement of the original models utilizing the discovered error slices, we leave results in Section A.7. Due to space limit, we put more examples including those of Spotlight and PlaneSpot in Section A.11, about 20 images for each identified slice.
@@ 5.1 Benchmark results: Dcbench
Dcbench (Eyuboglu et al. 2022) offers 886 publicly available settings for error slice discovery. Each setting consists of a trained ResNet-18 (He et al. 2016), a validation dataset and a test dataset, both with predefined underperforming slice labels. The validation dataset and its error slice labels are taken as input of slice discovery methods, while the test dataset and its error slice labels are used for evaluation. There are three types of slices in dcbench: correlation slices, rare slices, and noisy label slices. The correlation slices are generated from CelebA (Liu et al. 2015), while the other two types of slices are generated from ImageNet (Deng et al. 2009). More details are included in Section A.8. In terms of evaluation metrics, we employ precision@$k$ and average precision following dcbench’s practice, where precision@$k$ is the proportion of samples with top $k$ highest probabilities output by the learned slicing function that belongs to the predefined underperforming slice, and average precision is calculated based on precision@$k$ with different values of $k$. We also calculate manifold compactness as Equation 1. For all these metrics, a higher value indicates higher coherence of the identified slice, thus implying a more effective algorithm capable of error slice discovery.
Effectiveness of our method
Validity of our metric
It is also worth noting that the proposed metric manifold compactness shows a strong consistency with other metrics. Table 1 shows that the rank order of the four methods based on precision metrics is usually MCSD, Domino, Spotlight, PlaneSpot, the same as the rank order based on manifold compactness, except for the correlation slice where the rank order of Domino and Spotlight switches. While other metrics require access to predefined labels of underperforming slices, our metric does not. This demonstrates the validity and advantages of our proposed manifold compactness when measuring coherence and evaluating error slice discovery algorithms.
5.2 Case Study: CelebA
CelebA (Liu et al. 2015) is a large facial dataset of 202,599 images, each with annotations of 40 binary attributes. In the setting of subpopulation shift, it is the most widely adopted dataset since it is easy to generate spurious correlations between two specific attributes by downsampling the dataset (Yang et al. 2023; Sagawa et al. 2019; Liu et al. 2021a). Different from settings in dcbench, in this case study we follow Sagawa et al. (2019) to treat the binary label of blond hair as the target of prediction and directly use the whole dataset of CelebA (Liu et al. 2015) without downsampling, thus closer to the real scenario. In terms of implementation details, we employ the default data split provided by CelebA and follow the training process of ERM in (Sagawa et al. 2019) to train a ResNet-50. We apply error slice discovery algorithms on both categories respectively, thus taking advantage of outcome labels that are known during slice discovery. We also illustrate results of directly selecting top $\alpha n_{\text{te}}$ samples sorted by prediction losses.
@@ 5.2 Case Study: CelebA
[…]
@@
From Table 2, we can see that for both categories of CelebA, our algorithm identifies the most coherent underperforming slice in terms of manifold compactness, where higher is better. Although it ranks 2nd for the category of blond hair in terms of accuracy, where lower is better, for the task of error slice discovery, we put more emphasis on coherence since we want the identified slices to be interpretable, and we only require the performance of the slice to be lower than a threshold compared with the overall performance, as stated in Section 3. In Figure 4, left five columns and right five columns are from two categories separately. Four rows correspond to randomly sampled images from different sources: the error slice that Domino identifies, the error slice that MCSD identifies, top $\alpha n_{\text{te}}$ samples sorted by the loss, and all samples of the corresponding category. We can see that images from the error slice identified by MCSD obviously exhibit more coherent characteristics than others.
@@
[…]
@@ 5.3 Case Study: CheXpert
To demonstrate the effectiveness of our algorithm on other types of data, we conduct experiments on a medical imaging dataset, i.e. CheXpert (Irvin et al. 2019), where the task is to predict whether patients are ill or not based on their chest X-ray images. It contains 224,316 images from 65,240 patients. We follow the data split and training process of Yang et al. (2023) to train a ResNet-50. Still, we apply algorithms to images of ill and healthy patients respectively.
@@ 5.4 Case Study: BDD100K
[…]
@@
Compared with most previous algorithms (Eyuboglu et al. 2022; Wang et al. 2023; Plumb et al. 2023) requiring prediction probabilities as a part of input and are only designed for classification tasks, our algorithm is flexible to be employed in various tasks since it takes prediction losses as input. To illustrate its benefits of extending to other tasks, we conduct a case study on BDD100K (Yu et al. 2020), a large-scale dataset composed of driving scenes with abundant annotations. It includes ten tasks, of which we investigate object detection in our paper. The number of images in its object detection task is 79,863, which we split into train, validation, and test datasets with the ratio 2:1:1. We train a YOLOv7 (Wang, Bochkovskiy, and Liao 2023) and try to identify coherent error slices for it. We employ Average Precision (AP) as the performance metric that is widely adopted in detection tasks. Of the 13 categories in the task, we select 2 categories with a relatively high overall performance and a large sample size, i.e. pedestrian and traffic light. We apply our algorithm MCSD for each of them respectively. Note that we do not compare with Domino or PlaneSpot since neither of them is applicable to tasks other than classification.
Table 4 shows that MCSD successfully identifies error slices whose AP are lower than those of overall for both categories, and whose coherence is higher than that of Spotlight in terms of manifold compactness. In Figure 6, each row corresponds to five images randomly sampled from a given source. Left three columns correspond to the category of pedestrians, while right three columns correspond to the category of traffic lights. For both pedestrians and traffic lights, samples from the source of MCSD are coherent in that they are all taken at night. This conforms to the intuition that it is more difficult to recognize and locate objects when the light is poor. However, directly sampling from the high-loss images can hardly exhibit any common patterns. This reveals the potential of MCSD to extend to other tasks.
@@ 5.5 Case Study: CivilComments
In addition to experiments on visual tasks, to demonstrate the applicability of our method to other types of data, we conduct experiments on CivilComments (Borkan et al. 2019), a text dataset of 244,436 comments included in popular distribution shift benchmarks (Yang et al. 2023; Koh et al. 2021). Its task is to predict whether a given comment is toxic or not. We follow the data split and training process of Yang et al. (2023) to train a BERTbase. We apply algorithms to toxic and non-toxic comments respectively. In Table 5, we can see that MCSD identifies slices of the lowest accuracy and highest manifold compactness in both categories.
Table 5: Results on CivilComments, along with overall accuracy of the trained model. “Comp.” means “Manifold Compactness”. $\uparrow$” indicates that higher is better, while $\downarrow$” indicates that lower is better. We mark the best method in bold type. “%” indicates that the digits are percentage values.
@@
[…]
@@ 6 Conclusion
In this paper, inspired by the data geometry property, to better evaluate error slice discovery methods, we propose manifold compactness as a metric of slice coherence, which does not rely on predefined underperforming slice labels. Furthermore, with the help of explicit metrics for risk and coherence, we develop an algorithm that directly incorporates risk and coherence into the optimization objective, which is also flexible to be applied to different scenarios. We conduct experiments on a benchmark and perform multiple case studies to demonstrate both the validity of our proposed metric and the superiority of our algorithm, along with the potential to be flexibly extended to various types of tasks.
Acknowledgements
This work was supported by Tsinghua-Toyota Joint Research Fund, NSFC (No. 62425206, 62141607), and Beijing Municipal Science and Technology Project (No. Z241100004224009). Peng Cui is the corresponding author. All opinions in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Appendix
A.1 More results of preliminary studies of manifold compactness
@@
Comparison with Variance
We compare manifold compactness with variance, a common and representative metric of coherence directly calculated in the Euclidean space, on different semantically predefined slices. Note that since the slices are not of the same size, to compare manifold compactness of different slices properly, for each given slice we randomly sample a subset of size 150 with 20 times, and average the manifold compactness of 20 subsets as the manifold compactness of the given slice. For CelebA, we use the binary label $y$ to indicate whether the person has blond hair or not, and $a$ to indicate whether the person is male or not. From Table 6, we can see that for manifold compactness, its value of the more fine-grained slice, i.e. the more coherent slice, is larger than the more coarse-grained slice. For example, the manifold compactness of y=1&a=0y=1\&a=0 is $0.38$, larger than that of $y=1$ (the value is $0.35$) or $a=0$ (the value is $0.13$). Such a relationship holds for every pair of slices. However, in terms of variance, for example, variance of y=1&a=1y=1\&a=1 is $39.6$, larger than that of $y=1$ whose value is $36.2$, which is contrary to our expectation that variance of the more fine-grained slice is smaller than that of the more coarse-grained slice. For CheXpert, we use the binary label $y$ to indicate whether the person is ill or not, and $a$ to indicate whether the person is male or not. We also find that for manifold compactness, its value of the more fine-grained slice, i.e. the more coherent slice, is larger than the more coarse-grained slice, while the value of variance is not consistent with the granularity of the slice. We also additionally compare with other metrics that are directly calculated in Euclidean distance, including Mean Absolute Deviation (MeanAD), Median Absolute Deviation (MedianAD), and Interquartile Range (IQR). We find that they exhibit similar phenomenons to variance, i.e. the metric value of the more coarse-grained slice is sometime even smaller than that of the more fine-grained slice, which contradicts our expectation. Thus we demonstrate that manifold compactness is better at capturing semantic coherence than variance does.
@@ A.3 Theoretical analyses
We initialize $\mathbf{w}^{(0)}=\mathbf{w}^{\prime}$. Then we sweep a variable $i$ from $1$ to n1n-1. For each iteration with $1\leq j\leq n-1$, we generate a new weight vector $\mathbf{w}^{(j)}$ by the following process. We can prove that $\mathbf{w}^{(j)}$ is one solution of Equation 2 and $w_{r_{i}}^{(j)}\in\{0,1\},\forall 1\leq i\leq j$ by mathematical induction, which is already satisfied for $j=0$.
Firstly, we assign $w^{(j)}_{r_{i}}=w^{(j-1)}_{r_{i}}$ for $1\leq i\leq j-1$ and $j+2\leq i\leq n$ and denote $C=w^{(j-1)}_{r_{j}}+w^{(j-1)}_{r_{j+1}}\in[0,2]$.
@@
A.4 Experimental details and related analyses
Details of baselines are listed below:
  • Spotlight (d’Eon et al. 2022): It learns a point in the embedding space as the risky centroid, and chooses the closest points to the centroid as the error slice.
  • Domino (Eyuboglu et al. 2022): It develops an error-aware Gaussian mixture model (GMM) by incorporating predictions into the modeling process of GMM.
  • PlaneSpot (Plumb et al. 2023): It combines the prediction confidence and the reduced two-dimensional representation together as the input of a GMM.
Note that in all our experiments, we apply algorithms to the validation dataset $\{x_{i}^{\text{va}},y_{i}^{\text{va}}\}_{i=1}^{n_{\text{va}}}$ to obtain the slicing function $g_{\varphi}$, and then employ $g_{\varphi}$ on the test dataset $\{x_{i}^{\text{te}},y_{i}^{\text{te}}\}_{i=1}^{n_{\text{te}}}$ to acquire the prediction probability of each test sample belonging to the error slice. We choose the top $\alpha n_{\text{te}}$ samples from the test dataset sorted by the prediction probabilities as the error slice $\hat{\mathcal{S}}$, and calculate evaluation metrics based on it. As for our method that outputs the error slice of the validation dataset instead of a slicing function, to compare with other methods, we additionally train a binary MLP classifier $g_{\varphi}:\mathcal{Z}\mapsto[0,1]$ on top of embeddings, by treating samples in the error slice as positive examples, and treat the rest as negative ones. However, it is worth noting that our method is effective at error slice discovery without this additional slicing function, which is illustrated in Section A.4. Meanwhile, we explain the choice of the feature extractor $h_{fe}$ and conduct analyses in Section A.4.
Ablation of the slicing function
In Algorithm 1, after acquiring the desired samples, we additionally train an MLP as the slicing function. This is because the standard evaluation process of error slice discovery, an error slice discovery method is applied to validation data to obtain the slicing function, and then the slicing function is applied to test data to calculate evaluation metrics and conduct case analyses. Such practice is also adopted by dcbench (Eyuboglu et al. 2022), the benchmark of error slice discovery. Thus we follow this practice by additionally train a slicing function after we obtain the desired samples, so that we can compare fairly with previous methods. However, even without training this slicing function, our method can still produce meaningful results of case studies. We show examples randomly sampled from optimized results of Equation 2 in Figure 11 and 12. We can see that there is still high coherence in the identified slices.